OpenStudy (anonymous):

how do i solve for x-intercept in: y = -2 (x + 8) + 2?

OpenStudy (anonymous):

Perform the algebra, so you should get y=-2x-16+2, which simplifies to y=-2x-14. Then set y=0 and solve for x. 0=-2x-14

OpenStudy (anonymous):

Look for a common factor. 0=-2(x+7) 0=-2 or 0=x+7 0 does not equal 2, so your left with 0=x+7, which means x=-7. Your coordinate for the x-intercept would be (-7, 0).

OpenStudy (anonymous):

Does this make sense? Let me know if you need me to explain anything.